11.25) Two numbers have the ratio 3:4. If 4 is added to each of the numbers the resulting ratio is 4:5. Find the numbers.
11.27) A line segment of length 120 inches is divided into three parts whose lengths are proportional to 3,4,5. Find the lengths of the parts.
: 11.25) Two numbers have the ratio 3:4. If 4 is added to each of the numbers the resulting ratio is 4:5. Find the numbers.
11.27) A line segment of length 120 inches is divided into three parts whose lengths are proportional to 3,4,5. Find the lengths of the parts.

Answer by stanbon(19020)

11.25) Two numbers have the ratio 3:4. If 4 is added to each of the numbers the resulting ratio is 4:5. Find the numbers.
Ans: 3:4 is the same as 3x:4x
Adding 4 to each of the numbers you get:
(3x+4)/(4x+4) = 4/5
5(3x+4) = 4(4x+4)
15x+20 = 16x+16
x = 4
------------
The numbers are 3x = 12 and 4x = 16
=========================================

11.27) A line segment of length 120 inches is divided into three parts whose lengths are proportional to 3,4,5. Find the lengths of the parts.
Ans:
3x + 4x + 5x = 120
12x = 120
x = 10
3x = 30
4x = 40
5x = 50
================
Cheers,
Stan H.